OpenStudy (anonymous):

PLEASE HELP! MUST GRADUATE! Simplify f(x) =(3x^2-4x+1)/(3x-1) . Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f() is undefined. Make sure to show your work.

OpenStudy (anonymous):

f(1/3) on the last one.

OpenStudy (anonymous):

MUST GRADUATE! <3

OpenStudy (campbell_st):

factor the numerator (3x-1)(x-1) \[\frac{(3x-1)(x-1)}{(3x -1)}\] cancel the common factor of (3x -1) leaving (x -1) now using the simplified version (x -1) f(1) look at the factor (x -1) substituting x = 1 give 1 -1 = 0 next f(0) same process x = 0 gives 0 - 1 = -1 f(-1) same process x = -1 then -1 -1 = -2 the denominator of the fraction is (3x -1) then you need to solve 3x -1 = 0 x = 1/3 when x = 1/3 the denominator will have a value of zero which is undefined

OpenStudy (anonymous):

Good explanation, bro. The most important concept here is that you can't divide by 0. So if you plug in some x and get \[\frac{anything}{0}\] Then there's no answer and the y value for that x is undefined.